OpenStudy (richyw):

Need help with an integration. Even if someone knows what approach to use I should probably be OK

OpenStudy (anonymous):

what is the problem?

OpenStudy (zzr0ck3r):

seems to be very long

OpenStudy (richyw):

\[(2\pi a)^{-1/2} \int\limits_{-\infty}^{\infty}e^{2ik_ox}e^{-x^2/2a^2}dx\]

OpenStudy (richyw):

sorry my computer just froze

OpenStudy (unklerhaukus):

compair with the error function; \(\text{erf} (x)\)

OpenStudy (richyw):

ah ok. so wikipedia is explaning this to me. does this mean I need to use a taylor series expansion?

OpenStudy (unklerhaukus):

i dont think series expansion will help

OpenStudy (richyw):

oh no I see the antiderivative I need below that. I'll give it a shot now. Thanks a lot!

OpenStudy (unklerhaukus):

good luck